OpenStudy (batman31):

can I find an arc length in terms of Pi once i already have the sector area and my work?

OpenStudy (batman31):

if not, how do I find arc length?

zepdrix (zepdrix):

What are you asking Batman? :o

OpenStudy (batman31):

I'm saying I don't know how to find the length of XPY, I instead just managed to find the area of the sector, and i'm wondering how i would find the length. area=14.14 In terms of pi it is 4.5.

zepdrix (zepdrix):

Circumference of the circle is \(\Large\rm C=2\pi r\). But we only want to go 3/4 of the way around the circle. So, \(\Large\rm \frac{3}{4}C=\frac{3}{4}\cdot 2\pi r\)

zepdrix (zepdrix):

Don't think of area, think of circumference, distance around.

OpenStudy (batman31):

ok

OpenStudy (batman31):

12pi

OpenStudy (batman31):

that is the circumference, right?

zepdrix (zepdrix):

Mmm yes good. That's the total distance AROUND the circle. To find our arc length, we only want to go 3/4 of the distance around, yes?

zepdrix (zepdrix):

Notice that the small sector is a 90 degree angle. That accounts for 1/4 of the total pie.

OpenStudy (batman31):

0h. the answer is 9 pi. thank you good sir.

zepdrix (zepdrix):

yay good job \c:/